# QuickSplat: Fast 3D Surface Reconstruction via Learned Gaussian Initialization

### v1

@@ Abstract
Surface reconstruction is fundamental to computer vision and graphics, enabling applications in 3D modeling, mixed reality, robotics, and more. Existing approaches based on volumetric rendering obtain promising results, but optimize on a per-scene basis, resulting in a slow optimization that can struggle to model under-observed or textureless regions. We introduce QuickSplat, which learns data-driven priors to generate dense initializations for 2D gaussian splatting optimization of large-scale indoor scenes. This provides a strong starting point for the reconstruction, which accelerates the convergence of the optimization and improves the geometry of flat wall structures. We further learn to jointly estimate the densification and update of the scene parameters during each iteration; our proposed densifier network predicts new Gaussians based on the rendering gradients of existing ones, removing the needs of heuristics for densification. Extensive experiments on large-scale indoor scene reconstruction demonstrate the superiority of our data-driven optimization. Concretely, we accelerate runtime by 8x, while decreasing depth errors by up to 48% in comparison to state of the art methods.
@@ 1 Introduction
Recently, 3D Gaussian Splatting (3DGS) [27] achieves photorealistic novel-view-synthesis from multi-view images as input. It parameterizes a scene as a set of Gaussian primitives that are initialized from SfM and subsequently optimized and densified using reconstruction losses and a differentiable volumetric renderer. Subsequent works extend 3DGS to also obtain accurate surface reconstructions [25, 55, 22, 46]. However, these methods typically optimize each scene separately, i.e., many iterations of gradient descent are required, which can take over 30 minutes for large-scale indoor scenes [52]. Additionally, the surface is only optimized from the observed input images, but capturing sufficiently many diverse images remains challenging for large scenes. The resulting geometry may thus contain missing or deformed regions where there is less view coverage or texture information.
@@ 3.3 Iterative Gaussian Optimization
Optimizer Inspired by G3R [10], we learn an optimizer network $\theta_{O}$, that predicts updates for all voxels from $\{\mathcal{G},\nabla\mathcal{G}\}$ as input. In other words, we learn the function $f_{\theta_{O}}(\mathcal{G}_{t},\nabla\mathcal{G}_{t},t)=\Delta\mathcal{G}_{t}$, where $\Delta\mathcal{G}_{t}$ denotes the predicted update for all latent voxel features. Similar to G3R [10], we frame this process over multiple timesteps $t$, i.e., we iteratively calculate $\nabla\mathcal{G}_{t}$, predict $\Delta\mathcal{G}_{t}$, and update our representation as $\mathcal{G}_{t+1}{=}\mathcal{G}_{t}+\Delta\mathcal{G}_{t}$. We utilize a sparse 3D UNet architecture [12] for $\theta_{O}$. We normalize its output to lie within $[-1,1]$ to ensure that the updates do not overshoot.
@@
End-to-End Training We jointly train the densifier and optimizer networks in our second training stage. The initializer network $\theta_{I}$ remains frozen during this stage. We similarly calculate $\mathcal{L}_{\text{occ}}(\theta_{D})$ on the intermediate upsampling layers of the densifier. Additionally, we decode the updated voxel features $\mathcal{G}_{t+1}$ into Gaussians and calculate the rendering loss $\mathcal{L}(\theta_{O}){=}\mathcal{L}_{\text{c}}{+}\mathcal{L}_{\text{d}}{+}10\mathcal{L}_{\text{dist}}$. We run the densification-optimization loop for $T{=}5$ timesteps and calculate the losses after each timestep. Similarly to G3R [10], we detach the gradient of the losses for the subsequent timesteps, i.e., we optimize them separately.
@@ 4 Experiments
Figure 4: Qualitative comparison against baselines. We show top-down views of reconstructed mesh geometries (with and without vertex colors) in comparison to the ground-truth meshes from ScanNet++ [52]. Our method more accurately models flat wall structures and objects details, while producing fewer floating geometry artifacts.
@@
Dataset We train and evaluate our model on the ScanNet++ dataset [52]. We utilize 902 indoor scenes for training, where we filtered out some scenes with incomplete wall structures or very large extents. We train on the undistorted DSLR images at $360\times 540$ resolution and can double the resolution during evaluation to process images at $720\times 1080$. We evaluate our method on 20 unseen test scenes and report averaged metrics.
@@
Implementation Details We set $\text{v}_{\text{d}}{=}4$cm and predict $\text{v}_{\text{g}}{=}2$ Gaussians per voxel in all our experiments. The initializer and densifier network use four up/downsampling layers, and the optimizer UNet architecture follows G3R [10]. In total the networks have around 68M parameters. We set the learning rate to $1\mathrm{e}{-4}$ and train the networks for 3 days on a single Nvidia RTX A6000. In each iteration, we accumulate the gradients $\nabla\mathcal{G}$ from 100 training images. After running our iterative optimization for $t{=}5$ timesteps, we optionally refine the Gaussians for another 2000 steps of gradient descent (without adaptive density control). We denote results of our method as “w/ opt” that use this refinement and as “w/o opt” if they do not use it.
[…]
@@ 4.1 Comparison to State of the Art
Compared to SuGaR [22] and 2DGS [25] which are optimized per-scene using rendering losses, our method additionally leverages learned geometry priors from data. This is useful because indoor environments often have large textureless regions (e.g., white walls). Methods relying only on photometric errors may struggle in these areas (see Fig. 4 for examples). As a result, they can produce curved surfaces and floating artifacts. In contrast, our learned initializer network predicts a relatively dense Gaussian initialization, addressing the issue of missing SfM points in textureless regions.
Similar to ours, GS2Mesh [46] also utilizes geometry priors learned from data (i.e., the pretrained stereo estimator). However, since the synthesized views from 3DGS may not be as perfect as real images, there exists a noticeable domain gap for the stereo depth estimators when the renderings contain noise, generating high-frequency artifacts in the reconstructed geometry (see Tab. 1. In contrast, our method learns scene priors based on the Gaussian representation, and thus works better in this case.
Our method matches MonoSDF’s accuracy yet runs substantially faster. Moreover, unlike MonoSDF–which sometimes produces overly smooth surfaces and overlooks fine structures (e.g., the ladder in Fig. 5) when the monocular depth/normal estimator misses thin structures–our approach preserves more fine details.
@@ Appendix A More Implementation Details
     $\mathcal{\bar{G}}_{t}\leftarrow\mathcal{G}_{t}\cup\mathcal{\hat{G}}_{t}$triangleright Concatenate the new GS
@@
     $\mathcal{G}_{t+1}\leftarrow\mathcal{\bar{G}}_{t}+\Delta\mathcal{\bar{G}}_{t}$triangleright Update the parameters
@@
Appendix B More results
Figure 8: QuickSplat on ARKitscenes dataset without additional fine-tune. To show the generalization ability of QuickSplat, we use the model trained on ScanNet++ on ARKitscenes [5] dataset.
@@ Appendix B More results
Generalization.
To show the generalization ability of the method, we run QuickSplat trained on ScanNet++ directly to another indoor dataset: ARKitScenes [5] without additional fine-tuning. We process the ARKitScenes dataset following ScanNet++ to get the SfM point clouds and the alignment between the poses and GT mesh. The setting is more challenging due to the domain gap between the two datasets. In addition, the RGB captures in ARKitScenes have smaller FoV than ScanNet++’s, which also makes the task reconstructing from images more difficult. The results are shown in Fig. 8, which demonstrates the generalization ability of our proposed method.

### v2

@@ Abstract
Surface reconstruction is fundamental to computer vision and graphics, enabling applications in 3D modeling, mixed reality, robotics, and more. Existing approaches based on volumetric rendering obtain promising results, but optimize on a per-scene basis, resulting in a slow optimization that can struggle to model under-observed or textureless regions. We introduce QuickSplat, which learns data-driven priors to generate dense initializations for 2D gaussian splatting optimization of large-scale indoor scenes. This provides a strong starting point for the reconstruction, which accelerates the convergence of the optimization and improves the geometry of flat wall structures. We further learn to jointly estimate the densification and update of the scene parameters during each iteration; our proposed densifier network predicts new Gaussians based on the rendering gradients of existing ones, removing the needs of heuristics for densification. Extensive experiments on large-scale indoor scene reconstruction demonstrate the superiority of our data-driven optimization. Concretely, we accelerate runtime by 8x, while decreasing depth errors by 48% in comparison to state of the art methods.
@@ 1 Introduction
Recently, 3D Gaussian Splatting (3DGS) [28] achieves photorealistic novel-view-synthesis from multi-view images as input. It parameterizes a scene as a set of Gaussian primitives that are initialized from SfM and subsequently optimized and densified using reconstruction losses and a differentiable volumetric renderer. Subsequent works extend 3DGS to also obtain accurate surface reconstructions [26, 58, 23, 48, 10]. However, these methods typically optimize each scene separately, i.e., many iterations of gradient descent are required, which can take over 30 minutes for large-scale indoor scenes [55]. Additionally, the surface is only optimized from the observed input images, but capturing sufficiently many diverse images remains challenging for large scenes. The resulting geometry may thus contain missing or deformed regions where there is less view coverage or texture information.
@@ 3.3 Iterative Gaussian Optimization
Optimizer Inspired by G3R [11], we learn an optimizer network $\theta_{O}$, that predicts updates for all voxels from $\{\mathcal{G},\nabla\mathcal{G}\}$ as input. In other words, we learn the function $f_{\theta_{O}}(\mathcal{G}_{t},\nabla\mathcal{G}_{t},t)=\Delta\mathcal{G}_{t}$, where $\Delta\mathcal{G}_{t}$ denotes the predicted update for all latent voxel features. Similar to G3R [11], we frame this process over multiple timesteps $t$, i.e., we iteratively calculate $\nabla\mathcal{G}_{t}$, predict $\Delta\mathcal{G}_{t}$, and update our representation as $\mathcal{G}_{t+1}{=}\mathcal{G}_{t}+\Delta\mathcal{G}_{t}$. We utilize a sparse 3D UNet architecture [13] for $\theta_{O}$. We normalize its output to lie within [1,1][-1,1] to ensure that the updates do not overshoot.
@@
End-to-End Training We jointly train the densifier and optimizer networks in our second training stage. The initializer network $\theta_{I}$ remains frozen during this stage. We similarly calculate $\mathcal{L}_{\text{occ}}(\theta_{D})$ on the intermediate upsampling layers of the densifier. Additionally, we decode the updated voxel features $\mathcal{G}_{t+1}$ into Gaussians and calculate the rendering loss $\mathcal{L}(\theta_{O}){=}\mathcal{L}_{\text{c}}{+}\mathcal{L}_{\text{d}}{+}10\mathcal{L}_{\text{dist}}$. We run the densification-optimization loop for T=5T{=}5 timesteps and calculate the losses after each timestep. Similarly to G3R [11], we detach the gradient of the losses for the subsequent timesteps, i.e., we optimize them separately.
@@ 4 Experiments
Figure 4: Qualitative comparison against baselines. We show top-down views of reconstructed mesh geometries (with and without vertex colors) in comparison to the ground-truth meshes of ScanNet++ [55]. Our method more accurately models flat wall structures and objects details, while producing fewer floating geometry artifacts.
@@
Dataset We train and evaluate our model on the ScanNet++ dataset [55]. We utilize 902 indoor scenes for training, after filtering out some scenes with incomplete wall structures or very large bounding extents. We train on the undistorted DSLR images at $360\times 540$ resolution and double the resolution during evaluation to process images at $720\times 1080$. We evaluate our method on 20 unseen test scenes and report averaged metrics.
@@
Implementation Details We set $\text{v}_{\text{d}}{=}4$cm and predict $\text{v}_{\text{g}}{=}2$ Gaussians per voxel in all our experiments. The initializer and densifier network use four up/downsampling layers, and the optimizer UNet architecture follows G3R [11]. In total the networks have around 68M parameters. We set the learning rate to $1\mathrm{e}{-4}$ and train the networks for 3 days on a single Nvidia RTX A6000. In each iteration, we accumulate the gradients $\nabla\mathcal{G}$ from 100 training images. After running our iterative optimization for t=5t{=}5 timesteps, we optionally refine the Gaussians for another 2000 steps of gradient descent (without adaptive density control). We denote results of our method as “w/ opt” that use this refinement and as “w/o opt” if they do not use it.
Baselines We compare our method with several recent 3D surface reconstruction approaches: SuGaR [23], 2DGS [26], GS2Mesh [48], PGSR [10], and MonoSDF [57]. SuGaR optimizes 3DGS [28] and regularizes the Gaussians to align with surfaces, followed by mesh extraction using Poisson surface reconstruction. Similarly, GS2Mesh also begins by optimizing 3DGS, then renders stereo pairs and predicts depth maps using an off-the-shelf stereo depth estimator (i.e., DLNR [61]). 2DGS optimizes and densifies flat 2D Gaussians using gradient descent in a per-scene optimization lasting 30K iterations. PGSR renders unbiased depth maps from flattened 3D Gaussians and introduces both single-view and multi-view regularization losses to improve geometric reconstruction. MonoSDF leverages monocular depth and normal priors from a pre-trained model [17] to optimize the signed distance field (SDF) using differentiable volumetric rendering.
@@ 4.1 Comparison to State of the Art
Compared to SuGaR [23], 2DGS [26], and PGSR [10], which are optimized per scene using rendering losses, our method additionally leverages learned geometry priors from data. This is particularly useful in indoor environments, which often contain large textureless regions (e.g., white walls). Methods that rely solely on photometric errors may struggle in these areas (see Fig. 4 for examples), and therefore produce curved surfaces and floating artifacts. In contrast, our learned initializer network predicts a relatively dense Gaussian initialization, effectively addressing the issue of missing SfM points in textureless regions.
Similar to our approach, GS2Mesh [48] also utilizes geometry priors learned from data (i.e., from a pretrained stereo estimator). However, since synthesized views from 3DGS may not be as realistic as actual images, a noticeable domain gap exists for stereo depth estimators when the renderings contain noise. This can lead to high-frequency artifacts in the reconstructed geometry (see Tab. 1). In contrast, our method learns scene priors directly in the Gaussian representation space, allowing it to perform more robustly under such conditions.
Our method matches MonoSDF in accuracy while running substantially faster. Moreover, unlike MonoSDF—which sometimes produces overly smooth surfaces and misses fine structures (e.g., the ladder in Fig. 5) when the monocular depth or normal estimator fails to capture thin geometry—our approach better preserves fine details.
@@ Appendix A More Implementation Details
PGSR. (Chen et al. 2024)
We uses the official code and optimize the scenes for 30,000 iterations, with single view and multi-view regularization loss after 7,000 iterations. Exposure compensation is not used as ScanNet++ has fixed camera exposure.
@@
  $\mathcal{\bar{G}}_{t}\leftarrow\mathcal{G}_{t}\cup\mathcal{\hat{G}}_{t}$ $\triangleright$ Concatenate the new GS
@@
  $\mathcal{G}_{t+1}\leftarrow\mathcal{\bar{G}}_{t}+\Delta\mathcal{\bar{G}}_{t}$ $\triangleright$ Update the parameters
@@
Appendix B Additional Results
Generalization.
To demonstrate the generalization ability of our method, we run QuickSplat trained on ScanNet++ directly on other indoor datasets, such as ARKitScenes [5] and Mip-NeRF 360 [4], without any additional fine-tuning.
We process the ARKitScenes dataset following the same procedure as ScanNet++, obtaining the SfM point clouds and the alignment between camera poses and the ground-truth mesh. For Mip-NeRF 360 (Room), we restore the absolute scale of the official COLMAP point cloud and poses using a monocular metric depth estimator [52].
This cross-dataset setting is more challenging due to the domain gap between datasets. Additionally, the RGB captures in ARKitScenes and Mip-NeRF 360 have a smaller field of view compared to ScanNet++, making reconstruction from images more difficult. We compare QuickSplat with 2DGS in Tab. 4 and Fig. 7, which demonstrate the generalization capability of our proposed method. Additional reconstruction results are shown in Fig. 8.
Figure 7: Ours vs. 2DGS on ARKitScenes and MipNeRF 360. To demonstrate the generalization ability of QuickSplat, we run our model on ARKitScenes [5] and Mip-NeRF 360 [4] without fine-tuning. Compared to 2DGS, QuickSplat produces more complete geometry
[…]
2DGS 0.6978 0.3590 0.6015 1780s
Ours 0.1775 0.7698 0.4301 111s
Table 4: Evaluation on ARKitScenes (5 scenes, no fine-tuning).
Figure 8: More reconstruction result of QuickSplat on ARKitscenes dataset.
@@ Appendix B Additional Results
Appendix C Additional Ablations
Steps
We ablate the number of steps for the learned optimizer and post-optimization in Tab. 5. We observe that the depth error decreases gradually over the 5 optimization steps. Additional SGD optimization steps lead to a plateau and require more time. On the other hand, the Chamfer distance changes only marginally due to the good global geometry generated by our learned initialization.
$T=0$ $T=1$ $T=2$ $T=5$ SGD=1k SGD=2k
[…]
Time (s) 0.6 5.7 11 26 77 124
Table 5: Ablation over time steps.
Optimization and densification
We experiment with combining QuickSplat initialization with the original 2DGS optimization and densification, instead of using our optimization and densification networks, under comparable time constraints. As shown in Tab. 6, the learnable optimization and densification networks achieve better reconstruction in finer details (i.e., the accuracy metrics with small thresholds). Although the original SGD optimization and densification benefit from our initialization, our full method remains more efficient.
[…]
Ours 2DGS w/o densify 0.0692 0.4650 0.7211 0.1571 39s
Ours 2DGS w/ densify 0.0668 0.4796 0.7338 0.1486 39s
Ours Ours 0.0732 0.5263 0.7674 0.1461 26s
Table 6: Ablation on optimization and densification. We compare Quicksplat’s optimizer and densifier with original 2DGS optimizization (w/ and w/o densificaation) under similar time frame.
Extend initializer to other method
We demonstrate that our initializer can be easily integrated into other Gaussian splatting variants, such as SAGS [44]. Note that we modified SAGS to use 2D Gaussians instead of 3D Gaussians as the representation for reconstructing 3D surfaces. As shown in Tab. 7, SAGS with our initialization performs significantly better than with SfM initialization. Moreover, our full method, with the learned optimization and densification, reconstructs scenes more accurately and efficiently than SAGS’s original optimization and densification.
[…]
SfM SAGS w/o densify 0.1292 0.2781 0.5093 0.2879 429s
Ours SAGS w/o densify 0.0692 0.4724 0.7297 0.1633 253s
Ours SAGS w/ densify 0.0669 0.4825 0.7381 0.1625 276s
Ours Ours 0.0732 0.5263 0.7674 0.1461 26s
Table 7: Combined with SAGS [44]. We show that our initializer can be easily integrated into other methods, resulting in improved performance. In addition, our learned densification and optimization are faster and more accurate than SAGS under the same initialization. (Note that we modified SAGS to output 2D Gaussian splats for surface reconstruction.)